OpenStudy (anonymous):

solve the proportion. 10/a = 15/30

jimthompson5910 (jim_thompson5910):

10/a = 15/30 10*30 = 15a 300 = 15a 15a = 300 ... a = ???

OpenStudy (anonymous):

i got 0.05

OpenStudy (anonymous):

so a = 5 correct ?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

15a = 300 a = 300/15 a = ???

OpenStudy (anonymous):

a=20

OpenStudy (anonymous):

10/a= 15/30 30*15=300 300/15= 20 a=20 hope i helped

jimthompson5910 (jim_thompson5910):

a = 20 is correct

OpenStudy (anonymous):

thanks